A bag contains 14 identical balls, 4 of which are red, 5 black and 5 white. Six balls are drawn from the bag. find the probability that
(i) 3 are red, (ii) at least two are white.
1
Expert's answer
2020-07-27T18:48:45-0400

Solution :

14C6 = 3003 combinations of 14 balls taken 6 at a time.


i)P(3 are red)= (3 red out of 4 red balls) and (3 out of total 10 balls)

"\frac{4c3*10C3}{14C6}" ="\frac{480}{3003}" =0.159


ii)

P(at least two white)=1- [P(0 white)+P(1 white)]

that is P(X"\geq 2)" =1-[P(X=0)+P(x=1)]


= 1- ["\frac{9C6}{14C6}" + "\frac{5c1*9c5}{14C6}]"


=1-["\frac{84}{3003}" + "\frac{5*126}{3003}]"


=1-["\frac{714}{3003}]" = "\frac{2289}{3003}" =0.7622